OpenStudy (anonymous):

New Car Sales New car sales have increased at a roughly linear rate. Sales, in millions of vehicles, from 1992 to 2005, are given in the table below. Let x represent the number of years since 1990. Year Sales Year Sales 1992 12.8 1999 17.0 1993 13.9 2000 17.4 1994 15.0 2001 17.2 1995 14.7 2002 16.8 1996 15.1 2003 16.7 1997 15.2 2004 16.9 1998 15.6 2005 17.0 A. Find the equation of the least squares line and the coefficient of correaltion B. Find the equation of the least squares line using only the data for every other year starting with 1993, 1995, and so on. Find the coefficient of correlation. C. Compare your results for parts a and b. What do you find? Why do you think this happens?

OpenStudy (anonymous):

New Car Sales New car sales have increased at a roughly linear rate. Sales, in millions of vehicles, from 1992 to 2005, are given in the table below. Let x represent the number of years since 1990. Year Sales Year Sales 1992 12.8 1999 17.0 1993 13.9 2000 17.4 1994 15.0 2001 17.2 1995 14.7 2002 16.8 1996 15.1 2003 16.7 1997 15.2 2004 16.9 1998 15.6 2005 17.0 A. Find the equation of the least squares line and the coefficient of correaltion the equation of the least squares line is y=.16x+219 and the coefficient of correaltion is .16 y=.16x+b b=221.3-.16(1379) 92 221.3-220.64 92 b=219 y=.16x+219 let x represent years x=1 corresponding to the year 1992. let y represent sales value for x =92,92,94,95,96,97,98,99,100,101,102,103,104,105 value for y =12.8, 13.9, 15.0, 14.7, 15.1, 15.2, 15.6, 17.0, 17.4, 17.2, 16.8, 16.7, 16.9, 17.0 value for ∑x = 1,379 value for ∑y = 221.3 value for ∑xy = 21,866.1 value for ∑x2 = 136,059 value for ∑y2 = 3,524 m=n(∑xy)−(∑x )(∑x )→÷n(∑x2) − (∑x)2 <<formula for m 92(21,866.1)-(1379)(221.3) 92(136,059)-(1,379)2 <<sub from the table 92(21,866.1)-(1379)(221.3) 92(136,059)-(1,379)2 <<multiply 2011681.2-305172.7 12517428-1901641 <<subtract 1706508.5 10615787 m=.16

OpenStudy (anonymous):

is it right?